Raj Krishnan bought a Honda Civic for $17,345. He put down $6,000 and financed the rest through the dealer at an APR of 7.95 percent for four years. What is the effective annual interest rate (EAR) if the loan payments are made monthly? (Round answer to 0 decimal places, e.g. 15%.) Effective Annual interest rate %
Answer #1

Effective annual rate = [(1+Interest rate for period)^n]-1

Interest rate annual = 7.95% = 0.0795

N = number of periods in a year = 12[months]

Interest rate for period = Interest rate annual/n =0.0795/12 = 0.006625

Effective annual rate =[(1+0.006625)^12]-1 = 1.08246-1 = 0.08246 = 8.25%

Effective annual rate =8.25%